OpenStudy (annipuppi):

I need help with algebra 1 review fast please

OpenStudy (annipuppi):

So my teacher gave us 50 questions to reveiw for that standardized test we have next weekend and he said there would be some stuff we didn't know on it that we will be learning next week but he expected us to get 100% on it because we are allowed to get help from people or the internet. I think I have all the other answers right but I need help one a few. 12. how many real number solutions does the equation have? -7x^2+6x+3=0 14. What are the solutions of the system? y=x^2+3x-4 y=2x+2 16. simplify the expression. √3+4√3 a. 5√6 b. 5√3 c. 3√6 d. 3√3 21. graph the function and identify the vertical asymptote. y=(1/x-4)-2 22. find the sum or difference. [-9 -1 7] [-2 0 8] [_____]+[ ____] [ 0 7 2] [-4 5 -1] 27. simplify -(3ab^2)^-3 34. simplify the radical expression by rationalizing the denominator. 5/√30 35. Solve the equation identify any extraneous solutions √a=-2 36. Simplify the rational expression. State any excluded values. 4x-4/x-1 37. divide (25x^2-40x+18/(5x-3) 38. suppose that y varies inversly with x any y=2 when x=6 what is the equation for the inverse variation?

OpenStudy (annipuppi):

 This is how #22 is supposed to look